# Sampling as Bandits: Evaluation-Efficient Design for Black-Box Densities

### v1

@@ 5.2 Lorenz Weather Model
Wilks (2005) proposed a modification of the celebrated Lorenz dynamical system for weather forecast (Lorenz, 1995). They assumed that the weather dynamical system has ‘slow’ and ‘fast’ variables, where the former is observable at weather stations and the latter is unobservable. The modified Lorenz model includes a stochastic noise that represents the influence of the unobservable variables. It consists of 40 observable variables $(x_{1}(t),\dots,x_{40}(t))\in\mathbb{R}^{40}$ dependent on time $t$, whose dynamics is described by the following differential equation
@@
where $\theta:=(\theta_{1},\theta_{2})\in\mathbb{R}^{2}$ is the parameter that determines the system, and $\eta_{k}(t)$ is the stochastic noise that shall be defined below. Here, the index $k$ of each variable is an integer modulo 40, meaning that k1=40k-1=40 and k2=39k-2=39 when $k=1$, and that $k+1=1$ when $k=40$. We suppose that the model is defined over the time interval $[0,4]$.
We consider a setting where the initial state of the 40 variables at time $t=0$ is known and the subsequent state of the 40 variables is observed for 20 days on a daily basis. One day corresponds to the time duration $0.2$ in this model. Hence, we observe a sample path of the 40 variables at 20 time indices equispaced in the interval $[0,4]$, starting from the known initial state at $t=0$. To simulate sample paths of the model, we solve the differential equation (14) by the 4th-order Runge-Kutta method with a small time step $\delta t=0.025$, as in Thomas et al. (2022). Wilks (2005) assumed that the stochastic noise, for each $k$, has the following first-order autoregressive structure
$\eta_{k}(t+\delta t)=\phi\eta_{k}(t)+\sqrt{1-\phi^{2}}\epsilon_{k}(t)$ (15)
where $\phi\in[0,1]$ is set to $0.4$ throughout, and $\epsilon_{k}(t)$ is a standard normal noise drawn independently at each time $t$. The initial value of the noise $\eta_{k}(0)$ is given by $\sqrt{1-\phi^{2}}\epsilon_{k}(0)$. For our experiment, we sampled the initial state of the 40 variables from a standard normal distribution. We then simulated the sample path observed at the 20 time indices under the parameter $(\theta_{1},\theta_{2})=(2.0,0.1)$.
@@
Figure 5: Visualisation of 100 BIS samples for the Lorentz weather model. The left panel is the contour colourmap of the target posterior. The middle panel shows the locations of the initial 10 random points (star) and the following 90 points selected by BIS (circle). The right panel shows the contour colourmap of the KDE of the 100 weighted samples.
Hakkarainen et al. (2012) suggested to use a synthetic likelihood to perform Bayesian inference of the Lorenz weather model; see e.g. Frazier et al. (2023) for introduction of synthetic likelihoods. Section˜D.2 details the summary statistics used for the synthetic likelihood in this experiment. Our aim is to sample from the posterior defined by the synthetic likelihood and a prior. We specified a uniform prior on the parameter domain $\Theta=[0.0,5.0]\times[0.0,0.5]$, following Järvenpää et al. (2021). At each $\theta$, the synthetic likelihood is computed using 10,000 sample paths that are sufficiently many so that the estimation error of the synthetic likelihood is negligible. For BIS, we used the same setting in Section˜5.1. Figure˜5 shows 100 samples obtained by BIS, together with a kernel density estimator (KDE) based on the weighted samples. For comparison, we applied standard importance sampling to the target posterior, where we used the 10,000 scaled Halton sequence as samples and assigned self-normalised importance weights to them. Figure˜5 visualises the target posterior, using the KDE based on that 10,000 weighted samples as a sufficiently accurate representation of the target density. Figure˜5 illustrates that the weighted samples of BIS, visualised by the KDE, recover the target posterior well. To verify that, we evaluated the MMD between the target posterior and the 100 weighted samples of BIS, where the reported value of the MMD was $0.008$.
@@ 5.3 G-and-K Model
The g-and-k model is an expressive parametric family that can model the location, scale, shape, and skewness of data. Notable applications of the g-and-k model include modelling of foreign exchange return data that can exhibit heavy-tail and skewed behaviours (Prangle, 2020). The g-and-k model on $\mathbb{R}$ is parametrised by a four-dimensional parameter $\theta=(\theta_{1},\theta_{2},\theta_{3},\theta_{4})\in\mathbb{R}^{4}$. Only a quantile function $Q(u\mid\theta)$ of the model is analytical, where
@@
for $u\in(0,1)$. Here $c$ is a hyperparameter that we set to $0.8$.
@@
Denote by $Q^{\prime}(u\mid\theta)$ the derivative of the quantile function with respect to $u$. Denote by $Q(x\mid\theta)^{-1}$ the inverse of the quantile function $x=Q(u\mid\theta)$, that is, $u=Q(x\mid\theta)^{-1}$. Rayner and MacGillivray (2002) showed that the density of the g-and-k model is equal to
$p(x\mid\theta)=\frac{1}{Q^{\prime}(Q(x\mid\theta)^{-1}\mid\theta)}.$ (17)
However, this density is not available in closed form because the inverse $Q(x\mid\theta)^{-1}$ can be only numerically available. This makes the evaluation of the log-likelihood computationally expensive, as it requires numerical optimisation to compute the inverse $Q(x\mid\theta)^{-1}$ at each $x$ and $\theta$. ABC is arguably the most common approach to approximate Bayesian inference of the g-and-k model. Samples drawn by ABC follow an approximation of the exact posterior, involving a drawback to overestimate the variance of the posterior. In contrast, BIS performs exact Bayesian inference, targeting the exact posterior.
We simulated 1000 data points $\{x_{i}\}_{i=1}^{1000}$ from the g-and-k model $p(x\mid\theta_{0})$ under the parameter $\theta_{0}=(3,1,2,0.5)$. The domain $\Theta$ is the hyperrectangle $0,10]^{4}$ and a uniform prior on $\Theta$ is used to define the posterior. By BIS, we obtained 400 weighted samples of the posterior, where the first 40 points were randomly selected from the Halton sequence to increase stability of the GP surrogate model. In high dimensions, modes of probability measures with no sufficient tail decay can concentrate around the boundary of the domain, as known in the concentration on the sphere phenomenon. In this experiment, the mean function of the GP prior in GP-UJB was set to a quadratic polynomial whose coefficients are the hyperparameter, so that a sufficient tail decay is encoded in the surrogate density. The covariance kernel was set to the Gaussian kernel with length-scale $1$ and variance $20$, where the large variance assists in driving exploration of GP-UJB and avoiding overfitting of the surrogate density. The size of the candidate pool $M$ was fixed to $M=320,000$. Figure˜6 shows a histogram of the 400 samples of BIS in each coordinate of the parameter. Figure˜6 further shows a KDE constructed from the 400 weighted samples of BIS. BIS was compared with the KDE constructed from the $320,000$ scaled Halton sequence with self-normalised importance weights assigned. It demonstrates that the weighted samples of BIS recovers the target posterior, placing a probability mass correctly around the true parameter $\theta_{0}$.
Figure 6: Visualisation of 400 weighted samples of BIS for the g-and-k model. The top panels are the histograms of the samples of BIS in each coordinate of the parameter $\theta=(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$. The middle panels show the KDEs based on the weighted samples of BIS in each coordinate. The bottom panels show the marginal posteriors in each coordinate estimated by the KDEs based on the 320,000 weighted samples. The dotted line indicates the location of the correct parameter.
@@ 6 Related Work
To the best of the authors’ knowledge, BIS is the first method to frame the design of samples in sampling as a bandit problem. This work leveraged GP surrogate modelling of the target density to guide the design of samples adaptively. BIS allows for the convergence of the designed samples. This is in contrast to most existing works on experimental designs for GPs, in which the designed points are used only as training inputs of GP surrogate models.
Sinsbeck and Nowak (2017) proposed a GP surrogate model of the likelihood for efficient approximation of the posterior. They developed a sequential experimental design of points, at which the likelihood is evaluated for training of the GP surrogate model. Their approach is to find an estimator of the log-likelihood that minimises a Bayes risk associated with the GP surrogate model at each iteration. They then select a point that maximises the Bayes estimator of the likelihood. Once the specified number of points has been obtained, the posterior is approximated by a so-called “plug-in” posterior, in which the log-likelihood is replaced by the GP posterior mean. Järvenpää et al. (2021) extended the approach of Sinsbeck and Nowak (2017), considering a GP surrogate model of the posterior directly. They proposed a batched sequential design that selects a few points simultaneously at each iteration. However, their approaches draw on several heuristics and do not guarantee convergence of the GP surrogate model theoretically.
Recently, Kim and Sanz-Alonso (2024) proposed a simple modification of BO. Their approach is to include an extra random point in the training dataset of GP at each iteration. They applied the modified BO for maximum-a-posteriori estimation of the posterior, setting the objective function to the logarithm of the posterior. They showed that the mean function of the GP posterior used in the modified BO converges to the target posterior if the GP prior is well specified. That result is natural because adding random samples to the training dataset ensures that the target posterior is eventually evaluated everywhere in the domain. Nonetheless, the result offers an inspiring implication that imposing a certain restriction, by which points acquired by BO eventually fill the domain of interest, is sufficient to guarantee convergence of the GP surrogate model to the posterior.
Applications of GP surrogate modelling in Bayesian computation are not limited to approximation of the posterior density. Osborne et al. (2012) and Gunter et al. (2014) proposed a GP-based active learning framework to compute the model evidence. Meeds and Welling (2014) accelerated approximation Bayesian computation (ABC) with GP modelling of summary statistics. Gutmann and Corander (2016) developed an experimental design for GP modelling of discrepancy measures in ABC. Nemeth and Sherlock (2018) leveraged GP modelling to merge multiple posteriors conditional on different subsets of data. Bai et al. (2024) proposed a physics-informed GP surrogate model of forward simulators in Bayesian inverse problems.
@@ 7 Conclusion
Sampling from black-box densities has increasingly been a common task, for example, in Bayesian inference of complex models. However, there has been little existing literature on particle-based sampling methods tailored for black-box densities. This paper introduced BIS, the first importance sampling approach to black-box densities, which frames the design of samples as a bandit problem.
BIS is a simple and powerful sampling framework built on a choice of the proposal sequence and the point-selection criterion. This paper introduced our default point-selection criterion, GP-UJB, based on GP surrogate models of target densities. GP-UJB admits a decomposition into two terms that drive exploration of new points and exploitation of past information. On a theoretical level, we established the weak convergence of the weighted samples obtained by BIS. The empirical performance of BIS was demonstrated using benchmark densities and real-world models.
In the general framework of BIS, a choice of the point-selection criterion is flexible, whereas we focused on GP-UJB in this paper. As mentioned in Section˜3, difficulty in applying GPs in high dimensions is a well-known open challenge. GP-UJB inherits the same challenge for BIS. Designing point-selection criteria suitable for high-dimensional densities requires further substantial studies. Promising directions of future research include such new designs of point-selection criteria beyond the scope of GP surrogate modelling.
Appendix
This appendix contains details of the results presented in the main text. Appendix˜A presents preliminary lemmas useful for subsequent proofs. Appendix˜B provides proofs of all the theoretical results presented in the main text. Appendix˜C delivers simulation studies on the settings of BIS. Appendix˜D describes additional details of the experiments. It is convenient to introduce the following multi-index notations that will be used throughout.
@@
Let $\alpha=(\alpha_{1},\dots,\alpha_{d})$ be a $d$-dimensional vector of binary values, that is, $\alpha\in\{0,1\}^{d}$. We call $\alpha$ a binary multi-index. Let $|\alpha|$ be the number of non-zero elements of $\alpha$. Recall that $\partial_{i}$ denotes the partial derivative of a function with respect to the i-th coordinate of the argument. Let $\partial_{i}^{0}$ be an identity operator and let $\partial_{i}^{1}$ be the partial derivative $\partial_{i}$. Denote by $\partial^{\alpha}$ the mixed partial derivaitve $\partial^{\alpha}=\partial_{1}^{\alpha_{1}}\cdots\partial_{d}^{\alpha_{d}}$. Denote by $\theta^{\alpha}$ a $|\alpha|$-dimensional vector consisting of all i-th coordinates of $\theta$ such that $\alpha_{i}=1$. Denote by $\Theta^{\alpha}$ a set of all values of $\theta^{\alpha}$. Denote by $\theta_{*}^{\alpha}$ a $d$-dimensional vector s.t. the i-th coordinate is equal to that of $\theta$ if $\alpha_{i}=1$ or fixed to a constant $b_{i}$ otherwise, where $b_{i}$ denotes the right boundary of the i-th coordinate of the hyperrectangular domain $\Theta$.

### v2

@@ 6.2 Lorenz Weather Model
Wilks2005 proposed a stochastic modification of the Lorenz model for weather forecast (Lorenz1995). The formulation distinguishes between ‘slow’ observable and ‘fast’ unobservable variables, where the influence of the latter is modeled as a stochastic noise term. The system consists of 40 observable variables $(x_{1}(t),\dots,x_{40}(t))\in\mathbb{R}^{40}$, whose time evolution is described by
@@
where $\theta:=(\theta_{1},\theta_{2})\in\mathbb{R}^{2}$ represents the system parameter, and $\eta_{k}(t)$ denotes the stochastic noise defined subsequently. Here, the indices $k$ follow cyclic boundary conditions, meaning that $x_{0}=x_{40}$, $x_{-1}=x_{39}$, and $x_{41}=x_{1}$. The model is defined over the time interval $[0,4]$.
We assume the initial state of the 40 variables at $t=0$ is known, followed by daily observations over a 20-day period. In this model, one day corresponds to $0.2$ time units; thus, we observe a sample path at 20 equidistant time points within the interval $[0,4]$. Numerical integration of the differential equation (15) is performed using the fourth-order Runge-Kutta method with a time step of $\delta t=0.025$, following Thomas2022. Adopting the formulation of Wilks2005, the stochastic noise $\eta_{k}(t)$ is modeled as a first-order autoregressive process:
$\eta_{k}(t+\delta t)=\phi\eta_{k}(t)+\sqrt{1-\phi^{2}}\epsilon_{k}(t),$ (16)
where the autoregressive coefficient is fixed at $\phi=0.4$, and $\epsilon_{k}(t)$ represents i.i.d. standard normal noise. The initial noise term $\eta_{k}(0)$ is set to $\sqrt{1-\phi^{2}}\epsilon_{k}(0)$. In our experiments, the initial states of the variables were drawn from a standard normal distribution, and the observed sample path was simulated using the ground-truth parameters $(\theta_{1},\theta_{2})=(2.0,0.1)$.
@@
Figure 6: Visualization of 100 samples generated by BIS for the Lorenz model. The left panel is the contour colourmap of the target posterior. The middle panels show the samples, distinguishing between the initial 10 points (stars) and the subsequent 90 points selected by BIS (circles). The right panel shows the contour colourmap of the KDE of the 100 samples with the associated weights.
Hakkarainen2012 proposed using a synthetic likelihood to perform Bayesian inference of the Lorenz model; see Frazier2023 for a general introduction of synthetic likelihoods. The summary statistics defining the synthetic likelihood are detailed in Section˜D.2. We employed a uniform prior over the parameter domain $\Theta=[0.0,5.0]\times[0.0,0.5]$, following Jarvenpaa2021. To ensure that the estimation error of the synthetic likelihood is negligible, we computed it using 10,000 simulation paths at each query point $\theta$. The experimental configuration for BIS follows that of Section˜6.1. Figure˜6 presents the 100 samples generated by BIS, alongside a kernel density estimator (KDE) derived from the weighted samples. For comparison, we constructed a ground-truth proxy using standard importance sampling with 10,000 points from a scaled Halton sequence and self-normalized weights. The figure demonstrates that the KDE obtained by BIS accurately reconstructs the target posterior. This observation is quantitatively supported by the MMD between the target posterior and the 100 BIS samples, which was $0.008$.
@@ 6.3 G-and-K Model
The g-and-k model is an expressive parametric family capable of capturing location, scale, skewness, and kurtosis. Notable applications include modeling of foreign exchange returns, which frequently exhibit heavy tails and asymmetry (Prangle2020). The model is parameterized by a vector $\theta=(\theta_{1},\theta_{2},\theta_{3},\theta_{4})\in\mathbb{R}^{4}$. It is defined explicitly through its quantile function $Q(u\mid\theta)$, as the density is unavailable in closed form. For $u\in(0,1)$, the quantile function is given by
@@
where $z(u)$ denotes the standard normal quantile function, and $c$ is a constant fixed at $0.8$.
Figure 7: Marginal distributions for the g-and-k model parameters $\theta=(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$. The top panels display the histograms of 400 samples generated by BIS without weights. The middle panels display the KDEs derived from the 400 samples with the associated importance weights. The bottom panels show ground-truth marginals. Vertical dotted lines indicate the true parameter values.
@@
Let $Q^{\prime}(u\mid\theta)$ denote the derivative of the quantile function with respect to $u$. Rayner2002 showed that the density $p(x\mid\theta)$ can be expressed in terms of the quantile function and its inverse, denoted as $u=Q^{-1}(x\mid\theta)$. This expression is given by
$p(x\mid\theta)=\frac{1}{Q^{\prime}(Q^{-1}(x\mid\theta)\mid\theta)}.$ (18)
However, evaluating this density is computationally expensive because the inverse $Q^{-1}(x\mid\theta)$ has no closed-form expression and must be computed numerically for each data point $x$. While approximate Bayesian computation (ABC) is a standard approach to bypass this intractability, it targets an approximation of the posterior, which often results in variance inflation. In contrast, BIS performs exact Bayesian inference by directly targeting the true posterior density.
We generated a synthetic dataset of 1,000 observations $\{x_{i}\}_{i=1}^{1000}$ from the g-and-k model with true parameters $\theta_{0}=(3,1,2,0.5)$. The parameter domain was defined as the hyperrectangle $\Theta=[0,10]^{4}$, with a uniform prior placed over this domain. BIS was configured to produce 400 weighted samples, initialized with 40 points from the scaled Halton sequence to stabilize the GP surrogate. To encode sufficient tail decay, we assigned a quadratic polynomial mean function with coefficient hyperparameters to the GP prior, mitigating the risk of probability mass escaping outward due to the concentration of measure. We employed a Gaussian kernel with length-scale and variance hyperparameters for the covariance. The candidate pool size was set to $M=320,000$. Figure˜7 displays the marginal histograms for each parameter coordinate derived from the 400 BIS samples, along with the corresponding KDE. For benchmarking, we compared these results against ground-truth marginals constructed from 320,000 samples of the scaled Halton sequence with self-normalized importance weights. The figure confirms that the BIS samples accurately reconstruct the target posterior, correctly concentrating probability mass around the true parameter $\theta_{0}$.
6.4 US Precipitation Anomalies
Our final experiment demonstrates the practical utility of BIS through a large-scale real-world application involving MRFs. The dataset, previously analyzed in Kaufman2008 and Bolin2024, consists of annual precipitation anomalies for the year 1962, observed at $N=7352$ weather stations across the United States (see Figure˜8). We employ a standard Gaussian MRF (Rue2005) to model this dataset. Notably, Lyne2015 highlighted this class of models as a potential failure case for their posterior sampling approach relying on stochastic approximation of the computationally expensive likelihood. This challenge motivates the use of BIS, which evaluates the exact likelihood while minimizing the computational burden through efficient evaluation.
Figure 8: Visualization of the United States annual precipitation anomalies in 1962.
Let $\mathbf{y}=(y_{1},\dots,y_{N})\in\mathbb{R}^{N}$ denote the observations. Let $\mathbf{X}$ be a latent Gaussian MRF defined on a fixed triangulation of the continental United States. The covariance structure of the MRF is governed by the Matérn kernel with range and variance parameters $\rho$ and $\sigma$, respectively. The smoothness parameter of the Matérn kernel is fixed at $2.0$ to ensure that the latent field is at least once differentiable. Let $\mathbf{Q}(\rho,\sigma)$ be the precision matrix of the stochastic partial differential equation (SPDE) associated with the Matérn kernel. See, e.g., Lindgren2011 for the SPDE formulation of the Matérn kernel. Let $\mathbf{A}$ be the projection matrix mapping the latent field $\mathbf{X}$ from the triangulation to the observation locations. This hierarchical model is then expressed as:
$\mathbf{y}\mid\mathbf{X},\rho,\sigma,\sigma_{0}\sim\mathcal{N}(\mathbf{A}\mathbf{X},\sigma_{0}^{-2}\mathbf{I});\qquad\mathbf{X}\mid\rho,\sigma\sim\mathcal{N}(\mathbf{0},\mathbf{Q}(\rho,\sigma)^{-1}),$ (19)
where $\sigma_{0}$ denotes the scale of the observation noise.
Our interest lies in Bayesian inference of the three-dimensional parameter vector $\theta=(\rho,\sigma,\sigma_{0})$. Since the posterior distribution of the latent field $\mathbf{X}$ is Gaussian, $\mathbf{X}$ can be marginalized out analytically. The marginal log-likelihood $l(\theta)$ is explicitly available as:
$2l(\theta)=$ $N\log\left(\sigma_{0}^{2}\right)+\log\det\left(\mathbf{Q}(\rho,\sigma)\right)-\log\det\left(\mathbf{Q}(\rho,\sigma)+\sigma_{0}^{2}\mathbf{A}^{\mathrm{T}}\mathbf{A}\right)$ (20)
$-\sigma_{0}^{2}\mathbf{y}^{\mathrm{T}}\mathbf{y}+\sigma_{0}^{4}\mathbf{y}^{\mathrm{T}}\mathbf{A}\left(\mathbf{Q}(\rho,\sigma)+\sigma_{0}^{2}\mathbf{A}^{\mathrm{T}}\mathbf{A}\right)^{-1}\mathbf{A}^{\mathrm{T}}\mathbf{y}-N\log\left(2\pi\right).$ (21)
Although the likelihood is available in closed form, its evaluation presents a significant computational challenge due to the determinant and inverse operations involving massive precision matrices. While the sparsity of $\mathbf{Q}(\rho,\sigma)$ mitigates the cost, these linear algebra operations scale superlinearly as the dataset size increases. This necessitates an evaluation-efficient strategy for posterior sampling. We selected this application because, while the precision matrix is high-dimensional, performing MCMC remains feasible, allowing us to establish a robust baseline for comparison. We used the efficient implementation of the likelihood provided by Bolin2020.
For Bayesian inference, we assigned a uniform prior to $\theta$ over the compact domain $\Theta:=[0.1,5.0]^{3}$. The interval for the range parameter $\rho$ corresponds to approximately 0%–10% of the diameter of the observation region. The same interval was used for the variance parameter $\sigma$ and noise scale $\sigma_{0}$, as it was large enough to cover 2 to 3 times the standard deviation of the observed values. BIS was configured to produce 200 weighted samples, initialized with 20 points from the scaled Halton sequence. To encode sufficient tail decay, we assigned a quadratic polynomial mean function with coefficient hyperparameters to the GP prior. We employed a Gaussian kernel with length-scale and variance hyperparameters for the covariance. The candidate pool size was set to $M=80,000$. For the baseline, we performed $10,000$ iterations of MCMC, with a burn-in period of $5,000$ iterations and thinning applied every $10$ iterations to the remaining $5,000$ samples. Figure˜9 presents the resulting posterior approximation, using the KDE of the MCMC samples as the ground truth. The proposed BIS algorithm achieves an accurate approximation of the target posterior using only $200$ samples. The ability of BIS to recover the posterior structure with a limited sample budget demonstrates its computational advantage.
Figure 9: Marginal distributions for the MRF parameters $\theta=(\rho,\sigma,\sigma_{0})$. The top panels display the histograms of 200 samples generated by BIS. The middle panels display the KDEs derived from the 200 samples with the associated weights. The bottom panels show baseline marginals by $MC$.
@@ 7 Related Work
To the best of our knowledge, BIS is the first framework to formulate the sample design in importance sampling as a bandit problem. In most existing experimental design approaches for GPs, selected points serve merely as training inputs for a surrogate model. In contrast, BIS leverages the GP surrogate to guide the adaptive design of samples that converge to the target density.
Sinsbeck2017 developed a sequential design strategy to efficiently approximate the likelihood via a GP surrogate. Their method utilizes the Bayes estimator of the log-likelihood that minimizes the Bayes risk associated with the GP surrogate, proposing the design that maximizes this Bayes estimator at each iteration. Once the specified number of iterations is finished, the posterior is approximated via a plug-in estimator, where the log-likelihood is replaced by the GP posterior mean. Jarvenpaa2021 extended this approach by modeling the posterior density directly and proposing a batched sequential design. However, these methods rely on several heuristics and lack theoretical guarantees regarding the convergence of the surrogate model to the true target.
More recently, Kim2024 proposed a modification of BO for posterior approximation. Their strategy augments the standard BO acquisition with an additional random point at each iteration to ensure exploration. They demonstrated that, under a well-specified GP prior, the mean function of the GP posterior converges to the log target posterior. While this result is expected, given that random sampling eventually covers the entire domain, it offers the important insight that imposing simple constraints to ensure domain coverage is sufficient to guarantee consistency of the GP surrogate.
Beyond density approximation, GP surrogate modeling has been widely adopted in Bayesian computation. Notable applications include computing model evidence via active learning (Osborne2012; Gunter2014), accelerating ABC through summary statistic modeling (Meeds2014) or discrepancy modeling (Gutmann2016), merging posteriors from data subsets (Nemeth2018), and acting as physics-informed surrogates for forward simulators in inverse problems (Bai2024).
@@ 8 Conclusion
Sampling from black-box densities has become an increasingly common task, particularly in Bayesian inference for complex models. Despite its importance, there is limited literature on particle-based sampling methods specifically tailored for such densities. To address this gap, we introduced BIS, the first importance sampling framework that formulates the sample design as a bandit problem.
BIS offers a simple yet powerful framework, defined by a proposal sequence and a point-selection criterion. In this work, we proposed GP-UJB as a default criterion, leveraging GP surrogate models of the target density. The GP-UJB criterion naturally decomposes into two terms that balance the exploration of new regions with the exploitation of existing information. Theoretically, we established the weak convergence of the weighted samples generated by BIS. Empirically, we demonstrated the efficacy of the method on both benchmark densities and real-world applications.
While this paper focused on GP-UJB, the general BIS framework allows for flexible choices of point-selection criteria. As aforementioned in Section˜1, applying GPs in high-dimensional spaces remains a well-known open challenge, a limitation inherited by GP-UJB. Consequently, designing criteria suitable for high-dimensional densities requires substantial further investigation. Promising directions for future research include developing novel point-selection strategies that extend beyond the scope of GP surrogate modeling.
Supplementary Material
This supplementary material contains details of the results presented in the main text. Supplement˜A presents preliminary lemmas useful for subsequent proofs. Supplement˜B provides proofs of all the theoretical results presented in the main text. Supplement˜C delivers simulation studies on the settings of BIS. Supplement˜D describes additional details of the experiments. It is convenient to introduce the following multi-index notations that will be used throughout.
@@
Let $\alpha=(\alpha_{1},\dots,\alpha_{d})$ be a $d$-dimensional vector of binary values, that is, $\alpha\in\{0,1\}^{d}$. We call $\alpha$ a binary multi-index. Let $|\alpha|$ be the number of non-zero elements of $\alpha$. Recall that $\partial_{i}$ denotes the partial derivative of a function with respect to the i-th coordinate of the argument. Let $\partial_{i}^{0}$ be an identity operator and let $\partial_{i}^{1}$ be the partial derivative $\partial_{i}$. Denote by $\partial^{\alpha}$ the mixed partial derivative $\partial^{\alpha}=\partial_{1}^{\alpha_{1}}\cdots\partial_{d}^{\alpha_{d}}$. Denote by $\theta^{\alpha}$ a $|\alpha|$-dimensional vector consisting of all i-th coordinates of $\theta$ such that $\alpha_{i}=1$. Denote by $\Theta^{\alpha}$ a set of all values of $\theta^{\alpha}$. Denote by $\theta_{*}^{\alpha}$ a $d$-dimensional vector s.t. the i-th coordinate is equal to that of $\theta$ if $\alpha_{i}=1$ or fixed to a constant $b_{i}$ otherwise, where $b_{i}$ denotes the right boundary of the i-th coordinate of the hyperrectangular domain $\Theta$.